# Quantum Estimation of the Stokes Vector Rotation for a General Polarimetric Transformation

### v2

@@ Abstract
Classical polarimetry is a well-established discipline with diverse applications across different branches of science. The burgeoning interest in leveraging quantum resources to achieve highly sensitive measurements has spurred researchers to elucidate the behavior of polarized light within a quantum mechanical framework, thereby fostering the development of a quantum theory of polarimetry. In this work, drawing inspiration from polarimetric investigations in biological tissues, we investigate the precision limits of polarization rotation angle estimation about a known rotation axis, in a quantum polarimetric process, comprising three distinct quantum channels. The rotation angle to be estimated is induced by the retarder channel on the Stokes vector of the probe state. The diattenuator and depolarizer channels, acting on the probe state, can be thought of as effective noise processes. We explore the precision constraints inherent in quantum polarimetry by evaluating the quantum Fisher information (QFI) for probe states of significance in quantum metrology, namely NOON, Kings of Quantumness, and Coherent states. The effects of the noise channels as well as their ordering is analyzed on the estimation error of the rotation angle to characterize practical and optimal quantum probe states for quantum polarimetry. Furthermore, we propose an experimental framework tailored for NOON state quantum polarimetry, aiming to bridge theoretical insights with empirical validation.
@@ I Introduction
Optical polarimetry is a field of study and a set of techniques concerning measurement and interpretation of the polarization information. Studying the polarization of light before and after it interacts (transmission, reflection, or scattering) with a medium, one can infer several optical and geometric properties of the sample. Based on the optical properties of the material under study different polarimetry techniques might be used, such as transmission polarimetry, ellipsometry. Mueller polarimeters comprehensively measure the change of polarization state of light providing 16 elements of the Mueller matrix of a sample [15]. The non-invasive nature and high precision of polarimetry makes it suitable for studying sensitive samples. Ellipsometry is widely used for many applications including measurement of the refractive index and thickness of thin films [3]. Polarimetry techniques can also be used to measure the polarization parameters of the biological samples. For instance thickness of the retinal nerve fiber layer (RNFL) can be used to diagnose glaucoma [12, 16, 17, 18, 19]. The basic principle is that, RNFL is birefringent, therefore it can be modeled by a linear retarder which rotates the Stokes vector of the incoming probe state. Estimation of this rotation angle yields an estimate of the thickness of the RNFL [17, 18, 19].
@@
Several works have been carried out investigating the ultimate sensitivity limit and possible improvement for measurement of the optical properties of samples. It is shown that the NOON states and Kings of Quantumness (anticoherent states) yield Heisenberg limit (HL) scaling for rotation sensing [42, 43, 25, 44, 45, 46, 47]. For measurement of optical losses in a bosonic channel it is shown that Fock and two-mode squeezed vacuum (TMSV) states are optimal, however their error scaling doesn’t surpass the standard quantum limit (SQL) [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Studies have shown that for sensing the depolarization parameter HL scaling is not possible. However, using correlated probes and ancilla-assisted schemes increases the estimation accuracy [60, 61, 62, 63]. In general, it was argued that no parameter which is encoded on the density matrix by a non-unitary operator or a convex combination of unitary operators can be estimated beyond shot noise limit [64]. At the same time, subsequent studies demonstrated that employing controlled and error-corrected methodologies in metrology and environmental monitoring could potentially reinstate the HL in specific cases [65, 66, 67, 68].
@@
Based on the framework developed in [38] and inspired by the studies in vision and applications of metrology in biological systems [8, 9, 10, 11, 12, 16, 17, 18, 19, 69, 70], we aim to assess the feasibility of using quantum polarimetry for estimation of rotation angle of stokes vector due to a birefringent medium, modeled by a linear retarder, in presence of diattenuation and depolarization. Studies on biological tissues have shown that, although the polar decomposition of Mueller matrix [71] in tissue polarimetry can yield reliable polarization properties in classical regime, this decomposition does not correspond to the underlying physical reality [9, 72, 73]. Furthermore, experimental evidence suggests that optical quantum entanglement can survive while photons are being transmitted through highly scattering biological tissues [74, 75, 76, 77]. Therefore, it is crucial to investigate the precision limits in quantum polarimetry considering non-commutativity of the elementary components of the Mueller matrix, which implies considering different composition orders of the quantum polarization channels.
@@
This manuscript is organized as follows. In Section II we introduce concepts in quantum polarimetry such as Stokes operators and their transformations due to polarization channels. The basic concepts in quantum metrology, i.e. quantum Fisher information (QFI) and quantum Cramer-Rao bound (QCRB) are introduced in Section III. In Section IV the results of our calculations for QFI using NOON, Kings of Quantumness [42, 43, 25, 44, 45, 46, 47] and coherent states as probe states are given. In Section V possible experimental implementation is discussed and finally, in Section VI we present our conclusions. In the appendix we present a brief explanation on Majorana representation of quantum polarization states and quantum concepts in optical polarization and introduce Kings of Quantumness and the concept of anticoherence.
@@ II Quantum Description of Polarized Light and Polarization Transformations
Figure 1: (a) A schematic representation of the general polarization channel. In this figure the order of the action of the diattunator, retarder and depolarizer channels follows the standard Li-Chipman form. (b) Distribution of fibers in an inhomogeneous biological sample and overall rotation axis experienced by the probe. The large gray arrow represents the net rotation axis for the entire sample and the smaller colored arrows in dashed circles represents the local net rotation axis for the respective fiber distribution at each specific subsection of the sample.
@@
in which $\mathbf{n}=(\Theta,\Phi)$ is the rotation axis with and $\theta$ is the angle of rotation. In composite, anisotropic mediums such as biological tissues, retardation phenomena may arise from various factors such as reflection, refraction at interfaces, or propagation through the medium. Notably, these materials harbor distributed fibers, such as collagen fibers, exhibiting non-uniform distribution throughout the sample. While initial models often assumed uniformity in the refractive index difference between ordinary and extraordinary axes, and the uniform direction of the extraordinary axis across the sample, recent investigations have begun to challenge this simplification [9]. The orientation of the retarder axis aligns with the prevailing direction of these fibers within a specific volume of the sample [9]. However, during local probing of the sample, fiber orientations may locally fluctuate across different subsections, as illustrated in Fig. 1(b). Moreover, reconstructing three-dimensional images from various sections to ascertain fiber direction introduces inherent noise susceptibility, leading to inevitable uncertainties in inferred orientations [96, 97, 98, 99, 100]. Consequently, among the retarder’s three parameters, two primarily dictate the resultant fiber orientation—thus determining the rotation axis, albeit subject to uncertainty arising from the aforementioned factors—while the remaining parameter quantifies the degree of rotation, facilitating estimation of sample thickness. Our focus lies solely on accurately estimating the rotation angle.
@@
in which $p$ is the depolarization parameter and $p_{N}$ is the weight factor in each photon number subspace. To describe the action of this channel then what one can do to increase the depolarization factor $p$. However, this decomposition doesn’t always work for the case of quantum states and also, such a convex combination can only describe single photon states and classical states undergoing isotropic depolarization [41, 101, 40]. We can construct a quantum depolarization channel by a convex combination of different rotation operators [40, 41].
@@
In Eq. 16, $P(i)$ is the probability for the rotation operator $\hat{R}_{i}$ and the sum of all the probabilities must add up to one. Depolarization can be caused the interaction of light with an ensemble of atoms due to random rotations induced by the interaction on the photonic state [102, 103, 104]. The channel given in Eq. 16 is a simple model for such a physical process [40]. However, there are also models for depolarization of quantum states based on master equations [102, 103, 104], two of which are presented in Eq. 17 and Eq. 18.
@@ IV Results
In this section we numerically calculate the QFI for the rotation angle $\theta$ in order to determine its ultimate measurement precision. In all of our calculations the rotation angle to be estimated is $\theta=\pi/10$, which falls into the range of small angles for which the Kings of Quantumness (introduced in appendix) yield optimal sensitivity [45] for every value of $S$ considered in this study. The probe states have been considered for this task are the coherent state, NOON states and Kings of Quantumness.
@@ VI Conclusion
In conclusion, we studied the precision limits for single parameter estimation of the rotation angle of the Stokes vector in a polarimetric setup, within the framework of quantum polarimetry. We have incorporated the effects of depolarization and diattenuation channels within our numerical calculations and considered different orders of implementations of these channels. For the sake of completeness, we have utilized three different depolarization channels found in the literature in our calculations. The results show that, for our chosen sets of parameters and operators, QFI greatly diminishes upon implementation of the depolarization channel. For depolarization based on convex combination of rotations and the isotropic depolarization, there exists a minimal difference between forward and reverse decomposition in the final result for the QFI for weak depolarization and the difference is slightly more pronounced in the strong depolarization regime. However, for the anisotropic depolarization, after applying the depolarization channel, the reverse and forward decompositions yield different behavior with the average number of photons for Kings of Quantumness and the NOON state. The scaling of coherent state remains monotonic with the average number of photons and both decompositions yield the same result. However, for NOON state and Kings of Quantumness QFI becomes non-monotonic with the average photon number in the strong depolarization regime. As expected, implementing diattenuation decreases QFI for all cases. However, our results demonstrate that metrological power of the NOON state diminishes more than that of Kings of Quantumness such that for larger average photon numbers the latter yields a higher QFI. Therefore our results suggest that, although HL scaling is lost after applying the noise channels, one can potentially utilize low photon number NOON states to gain a quantum advantage in polarimetric retardation measurements on a lossy medium. Possible experimental procedure for carrying out such a task is proposed.
@@ Acknowledgements.
We gratefully acknowledge financial support from the Scientific and Technological Research Council of Türkiye (TÜBİTAK), grant No. 120F200. This work has been funded by the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 899580); the German Federal Ministry of Education and Research (FKZ 13N14877 and 13N15956); and the Cluster of Excellence “Balance of the Microverse” (EXC 2051-Project No. 390713860). V.B. thanks for funding of this work also through the ProChance-career program of the Friedrich Schiller University Jena.
@@ Appendix A Majorana Constellations and Quantumness of Polarization States
Using a method called the Majorana spin representation, one can construct an arbitrary pure state with angular momentum $S$ as a superposition of $2S$ spin-1/2 systems. Note that although the Hilbert space of $S$ spin-1/2 systems $\mathcal{H}\in\mathbb{C}^{2^{S}}$ is of dimension $2^{S}$, it’s symmetric subspace, is of dimension $12S+$ 1 [127, 128]. Majorana showed that there exists a bijection between the pure spin state given in Eq. 24 and roots of the polynomial $M(z)$, which is called the Majorana polynomial [129, 127].
@@
Due to the isomorphism that is established by Schwinger boson representation between spin systems and two-mode bosonic systems, one can use Majorana’s method to represent any pure polarization state with $n=2S$ photons as a collection of points on the surface a Riemann sphere (Poincare sphere within the context of polarization optics).
@@
The methodology described so far, explains how to represent a general pure polarization state as a constellation of points on the surface of a sphere and its connection to the original method by Majorana developed for quantum angular momentum states. However, it is constructive to briefly describe another approach based on coherent states to make connections between the constellations, quasi-probability distribution and different notions of quantumness of polarization states. It is convenient to express the Bloch coherent (a.k.a spin coherent, atomic coherent and $\mathfrak{su}(2)$ coherent) state using Dicke basis.
@@
In Eq. 31, $\hat{D}$ is the displacement operator on $\mathcal{S}^{2}$. Any pure state in Dicke basis can be written as Eq. 24. The stellar function (a.k.a Fock-Bargmann representation) of the state for $\mathcal{H}_{S}$ is defined using the overlap between the state Eq. 24 and the Bloch coherent state [133, 134].
@@
In Eq. 41, the DOP of a state is defined as the distance between the Q function of that state and the Q function for the maximally unpolarized state ($Q_{\text{unpol}}=1/4\pi$). The DOP defined in Eq. 41 is more suitable for quantum states of light compared to Eq. 5 because, as explained previously, the higher order multipoles play a more significant role in the polarization behavior of the quantum states. However, in most of the cases, one can gain substantial amount of information on polarization characteristics of the state only considering a finite number of its multipoles. A useful quantity which encodes the polarization information of a quantum state up to its $M$th order is the $M$th order cumulative multipolar distribution defined below [39, 89, 135].

### v3

@@ Abstract
Classical polarimetry is a well-established discipline with diverse applications across different branches of science. The burgeoning interest in leveraging quantum resources to achieve highly sensitive measurements has spurred researchers to elucidate the behavior of polarized light within a quantum mechanical framework, thereby fostering the development of a quantum theory of polarimetry. In this work, drawing inspiration from polarimetric investigations in biological tissues, we investigate the precision limits of polarization rotation angle estimation about a known rotation axis, in a quantum polarimetric process, comprising three distinct quantum channels. The rotation angle to be estimated is induced by the retarder channel on the Stokes vector of the probe state. The diattenuator and depolarizer channels, acting on the probe state, can be thought of as effective noise processes. We explore the precision constraints inherent in quantum polarimetry by evaluating the quantum Fisher information for probe states of significance in quantum metrology, namely NOON, Kings of Quantumness, and Coherent states. The effects of the noise channels as well as their ordering is analyzed on the estimation error of the rotation angle to characterize practical and optimal quantum probe states for quantum polarimetry. Furthermore, we propose an experimental framework tailored for NOON state quantum polarimetry, aiming to bridge theoretical insights with empirical validation.
@@ I Introduction
Optical polarimetry is a field of study and a set of techniques concerning measurement and interpretation of the polarization information. Studying the polarization of light before and after it interacts (transmission, reflection, or scattering) with a medium, one can infer several optical and geometric properties of the sample. Based on the optical properties of the material under study different polarimetry techniques might be used, such as transmission polarimetry or ellipsometry. Mueller polarimeters comprehensively measure the change of polarization state of light providing 16 elements of the Mueller matrix of a sample [15]. The non-invasive nature and high precision of polarimetry makes it suitable for studying sensitive samples. Ellipsometry is widely used for many applications including measurement of the refractive index and thickness of thin films [3]. Polarimetry techniques can also be used to measure the polarization parameters of the biological samples. For instance, the thickness of the retinal nerve fiber layer (RNFL) can be used to diagnose glaucoma [12, 16, 17, 18, 19]. The basic principle is that, RNFL is birefringent, therefore it can be modeled by a linear retarder which rotates the Stokes vector of the incoming probe state. Estimation of this rotation angle yields an estimate of the thickness of the RNFL [17, 18, 19].
@@
Several works have been carried out investigating the ultimate sensitivity limit and possible improvement for measurement of the optical properties of samples. It is shown that the NOON states and Kings of Quantumness (anticoherent states) yield Heisenberg limit (HL) scaling for rotation sensing [42, 43, 25, 44, 45, 46, 47]. Kings of Quantumness are states which have vanishing expectation values for their Stokes operators and they are called Kings or anticoherent states of order $t$, if all higher moments of the Stokes operators up to $t$ are isotropic. More information on these states is provided in the appendix. For measurement of optical losses in a bosonic channel it is shown that Fock and two-mode squeezed vacuum (TMSV) states are optimal, however their error scaling doesn’t surpass the standard quantum limit (SQL) [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Studies have shown that for sensing the depolarization parameter HL scaling is not possible. However, using correlated probes and ancilla-assisted schemes increases the estimation accuracy [60, 61, 62, 63]. In general, it was argued that no parameter which is encoded on the density matrix by a non-unitary operator or a convex combination of unitary operators can be estimated beyond shot noise limit [64]. At the same time, subsequent studies demonstrated that employing controlled and error-corrected methodologies in metrology and environmental monitoring could potentially reinstate the HL in specific cases [65, 66, 67, 68].
@@
Based on the framework developed in [38] and inspired by the studies in vision and applications of metrology in biological systems [8, 9, 10, 11, 12, 16, 17, 18, 19, 69, 70], we aim to assess the feasibility of using quantum polarimetry for estimation of rotation angle due to a birefringent medium, modeled by a linear retarder, in presence of diattenuation and depolarization. Studies on biological tissues have shown that, although the polar decomposition of Mueller matrix [71] in tissue polarimetry can yield reliable polarization properties in classical regime, this decomposition does not correspond to the underlying physical reality [9, 72, 73]. Furthermore, experimental evidence suggests that optical quantum entanglement can survive while photons are being transmitted through highly scattering biological tissues [74, 75, 76, 77]. Therefore, it is crucial to investigate the precision limits in quantum polarimetry considering non-commutativity of the elementary components of the Mueller matrix, which implies considering different composition orders of the quantum polarization channels.
@@
This manuscript is organized as follows. In Section II we introduce concepts in quantum polarimetry such as Stokes operators and their transformations due to polarization channels. The basic concepts in quantum metrology, i.e. quantum Fisher information (QFI) and quantum Cramer-Rao bound (QCRB) are introduced in Section III. In Section IV the results of our calculations for QFI using NOON, Kings of Quantumness [42, 43, 25, 44, 45, 46, 47] and coherent states as probe states are given. In Section V a possible experimental implementation is discussed and finally, in Section VI we present our conclusions. In the appendix we present a brief explanation on Majorana representation of quantum polarization states and quantum concepts in optical polarization and introduce Kings of Quantumness and the concept of anticoherence.
@@ II Quantum Description of Polarized Light and Polarization Transformations
Figure 1: (a) A schematic representation of the general polarization channel. In this figure the order of the action of the diattunator, retarder and depolarizer channels follows the standard Lu-Chipman form. (b) Distribution of fibers in an inhomogeneous biological sample and overall rotation axis experienced by the probe. The large gray arrow represents the net rotation axis for the entire sample and the smaller colored arrows in dashed circles represents the local net rotation axis for the respective fiber distribution at each specific subsection of the sample.
@@
in which $\mathbf{n}=(\Theta,\Phi)$ is the rotation axis with and $\theta$ is the angle of rotation. In composite, anisotropic media such as biological tissues, retardation phenomena may arise from various factors such as reflection, refraction at interfaces, or propagation through the medium. Notably, these materials harbor distributed fibers, such as collagen fibers, exhibiting non-uniform distribution throughout the sample. While initial models often assumed uniformity in the refractive index difference between ordinary and extraordinary axes, and the uniform direction of the extraordinary axis across the sample, recent investigations have begun to challenge this simplification [9]. The orientation of the retarder axis aligns with the prevailing direction of these fibers within a specific volume of the sample [9]. However, during local probing of the sample, fiber orientations may locally fluctuate across different subsections, as illustrated in Fig. 1(b). Moreover, reconstructing three-dimensional images from various sections to ascertain fiber direction introduces inherent noise susceptibility, leading to inevitable uncertainties in inferred orientations [96, 97, 98, 99, 100]. Consequently, among the retarder’s three parameters, two primarily dictate the resultant fiber orientation—thus determining the rotation axis, albeit subject to uncertainty arising from the aforementioned factors—while the remaining parameter quantifies the degree of rotation, facilitating estimation of sample thickness. Our focus lies solely on accurately estimating the rotation angle.
@@
in which $p$ is the depolarization parameter and $p_{N}$ is the weight factor in each photon number subspace. However, this decomposition doesn’t always work for the case of quantum states and also, such a convex combination can only describe single photon states and classical states undergoing isotropic depolarization [41, 40]. Within the context of quantum optics, polarization is $\mathfrak{su}(2)$-invariant and an unpolarized state is defined as a state which remains invariant under any linear polarization transformation. A valid quantum optical depolarization model must preserve the invariant subspaces of the optical field and the steady-state in each of them must be a diagonal state [101, 102, 103]. We can construct a quantum depolarization channel by a convex combination of different rotation operators [40, 41].
@@
In Eq. 16, $P(i)$ is the probability for the rotation operator $\hat{R}_{i}$ and the sum of all the probabilities must add up to one. Depolarization can be caused by the interaction of light with an ensemble of atoms due to random rotations induced by the interaction on the photonic state [101, 102, 103]. The channel given in Eq. 16 is a simple model for such a physical process [40]. However, there are also models for depolarization of quantum states based on master equations [101, 102, 103], two of which are presented in Eq. 17 and Eq. 18.
@@ IV Results
In this section we numerically calculate the QFI for the rotation angle $\theta$ in order to determine its ultimate measurement precision. In all of our calculations the rotation angle to be estimated is $\theta=\pi/10$, which falls into the range of small angles for which the Kings of Quantumness yield optimal sensitivity [45] for every value of $S$ considered in this study. The probe states that have been considered for this task are the coherent state, NOON states and Kings of Quantumness.
@@ IV.B Noisy Rotation Sensing with Depolarization and Diattenuation
So far we have compared the precision for estimation of the rotation angle between the three states under various settings. Although the QFI quantifies the ultimate precision bound, achieving the precision specified by the QFI might be experimentally challenging. Therefore it is instructive to briefly discuss the possible achievable precision limits under readily available experimental measurement operators. For this purpose we present the result for the simplest case of a probe state with a single photon ($\left\langle n\right\rangle=1$). Considering the POVMs corresponding to the Pauli operators $\hat{\sigma}_{x}$, $\hat{\sigma}_{y}$ and $\hat{\sigma}_{z}$, which can be easily implemented experimentally, the Fisher information for the NOON and coherent states become $2$ and $2/e$ respectively. Therefore, it is possible to extract the quantum advantage in this simple case using experimentally accessible measurement operators. Extension of this methodology to the higher excitation subspaces of the probe field is tedious but straightforward. Additionally, to gain maximal information, one should utilize symmetric informationally complete (SIC) POVMs to perform the measurements. However, experimental feasibility of such SIC-POVMs and complete characterization of precision based on experimentally easy to implement measurement operators is out of the scope of this work.
@@ VI Conclusion
In conclusion, we studied the precision limits for single parameter estimation of the rotation angle of the Stokes vector in a polarimetric setup, within the framework of quantum polarimetry. We have incorporated the effects of depolarization and diattenuation channels within our numerical calculations and considered different orders of implementations of these channels. For the sake of completeness, we have utilized three different depolarization channels found in the literature in our calculations. All three depolarization models are consistent with the notion of polarization in quantum optics. The results show that, for our chosen sets of parameters and operators, QFI greatly diminishes upon implementation of the depolarization channel. For depolarization based on convex combination of rotations and the isotropic depolarization, there exists a minimal difference between forward and reverse decomposition in the final result for the QFI for weak depolarization and the difference is slightly more pronounced in the strong depolarization regime. However, for the anisotropic depolarization, after applying the depolarization channel, the reverse and forward decompositions yield different behavior with the average number of photons for Kings of Quantumness and the NOON state. The scaling of coherent state remains monotonic with the average number of photons and both decompositions yield the same result. However, for NOON state and Kings of Quantumness QFI becomes non-monotonic with the average photon number in the strong depolarization regime. This non-monotonicity manifests itself in all three depolarization models used in our calculations and therefore strongly suggests that there exists a law of diminishing returns for extracting information about the rotation parameter with respect to the number of excitations in the probe field in the higher range of the depolarization regime considered in this study. As expected, implementing diattenuation decreases QFI for all cases. However, our results demonstrate that metrological power of the NOON state diminishes more than that of Kings of Quantumness such that for larger average photon numbers the latter yields a higher QFI. Therefore our results suggest that, although HL scaling is lost after applying the noise channels, one can potentially utilize low photon number NOON states to gain a quantum advantage in polarimetric retardation measurements on a lossy medium. Possible experimental procedure for carrying out such a task is proposed.
@@ Acknowledgements.
We gratefully acknowledge financial support from the Scientific and Technological Research Council of Türkiye (TÜBİTAK), grant No. 120F200. This work has been funded by the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 899580); the German Federal Ministry of Education and Research (FKZ 13N14877 and 13N15956); and the Cluster of Excellence “Balance of the Microverse” (EXC 2051-Project No. 390713860). V.B. thanks for funding of this work also through the ProChance-career program of the Friedrich Schiller University Jena. A.P. thanks Aaron Z. Goldberg from National Research Council of Canada for fruitful discussions.
@@ Appendix A Majorana Constellations and Quantumness of Polarization States
Using a method called the Majorana spin representation, one can construct an arbitrary pure state with angular momentum $S$ as a superposition of $2S$ spin-1/2 systems. Note that although the Hilbert space of $S$ spin-1/2 systems $\mathcal{H}\in\mathbb{C}^{2^{S}}$ is of dimension $2^{S}$, its symmetric subspace, is of dimension $12S+$ 1 [126, 127]. Majorana showed that there exists a bijection between the pure spin state given in Eq. 24 and roots of the polynomial $M(z)$, which is called the Majorana polynomial [128, 126].
@@
Due to the isomorphism that is established by Schwinger boson representation between spin systems and two-mode bosonic systems, one can use Majorana’s method to represent any pure polarization state with $n=2S$ photons as a collection of points on the surface of a Riemann sphere (Poincare sphere within the context of polarization optics).
@@
The methodology described so far, explains how to represent a general pure polarization state as a constellation of points on the surface of a sphere and its connection to the original method by Majorana developed for quantum angular momentum states. However, it is constructive to briefly describe another approach based on coherent states to make connections between the constellations, quasi-probability distribution and different notions of quantumness of polarization states. It is convenient to express the Bloch coherent (also known as spin coherent, atomic coherent and $\mathfrak{su}(2)$ coherent) state using Dicke basis.
@@
In Eq. 31, $\hat{D}$ is the displacement operator on $\mathcal{S}^{2}$. Any pure state in Dicke basis can be written as Eq. 24. The stellar function (also known as Fock-Bargmann representation) of the state for $\mathcal{H}_{S}$ is defined using the overlap between the state Eq. 24 and the Bloch coherent state [132, 133].
@@
In Eq. 41, the DOP of a state is defined as the distance between the Q function of that state and the Q function for the maximally unpolarized state ($Q_{\text{unpol}}=1/4\pi$). The DOP defined in Eq. 41 is more suitable for quantum states of light compared to Eq. 5 because, as explained previously, the higher order multipoles play a more significant role in the polarization behavior of the quantum states. However, in most of the cases, one can gain substantial amount of information on polarization characteristics of the state only considering a finite number of its multipoles. A useful quantity which encodes the polarization information of a quantum state up to its $M^{\text{th}}$ order is the $M^{\text{th}}$ order cumulative multipolar distribution defined below [39, 89, 134].